a merchant has coffee worth $60 a pound that she wishes to mix with 50 pounds of coffee worth $90 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $60 coffee should be used?
Answer by mananth(16946)

--------price ---------------- quantity
Coffee I 90 ---------------- 50 lbs
coffee II 60 ---------------- x lbs
Mixture 70 ---------------- 50 + x lbs

90*50+60 x=70(50 +x)
4500+60 x=3500 + 70 x
60 x + -70 x = 3500 - -4500
-10 x = -1000
/ -10
x = 100
100 lbs of 60 % coffee II